JoelTheBoss (joel_the_boss):

The Math Club is sponsoring a bake sale. If their goal is to raise at least $200, how many pies must they sell at $4.00 each in order to meet that goal? Write, solve, and graph an inequality that represents this situation.

JoelTheBoss (joel_the_boss):

@ganeshie8 Mind helping me? Trying to refresh on basic algebra before i head into trig. Math has never been my strong point. Thanks. :)

JoelTheBoss (joel_the_boss):

goal is 200- g pies 4.00 each- 4 number of pies sold- n \(p*n=g\) \(4*n=200\) \(\frac{4}{4}*n= \frac{200}{4}\) \(n=50\) is that correct? @vera_ewing @ganeshie8

OpenStudy (butterflydreamer):

weeell firstly, you want to note that we're dealing with an inequality :) "...goal is to raise AT LEAST $200"... so this basically means " \[\ge 200\] Then we know each pie costs $4 So if n = number of pies and we need to raise at least $200, the inequality should be : \[4n \ge 200 \]

OpenStudy (butterflydreamer):

but yes, you're correct in the way you solved it :) At least 50 pies are needed! \[4n \ge 200 \rightarrow n \ge 50 \]

JoelTheBoss (joel_the_boss):

Thank you very much. :) @butterflydreamer

OpenStudy (butterflydreamer):

nooo problemo ^_^!